# Hedging Hydrogen: Planning and Contracting Under Uncertainty for a Green Hydrogen Producer

### v2

@@ 1.3 Contributions
An energy market focused 2-stage stochastic model of an electrolytic hydrogen production plant in the context of a HPA, with co-optimisation of plant capacities, multiple electricity sourcing options, and the possibility of green hydrogen subsidies.
@@
Hedging energy re-sale (resale-enabled or no-resale formulations).
@@ 2.1 Overview
HPA’s in practice may have different conditions that are designed to reduce demand uncertainty, including but not limited to minimum and maximum yearly, monthly, daily, and hourly volumes. In this study, a HPA with fixed demand corresponds to a contract where the consumer is bound to follow precisely their original profile (i.e., all volumes down to hourly volumes are fixed, and no randomisation is included). Uncertainty can then be added by relaxing some of these rules, for example, by allowing monthly volumes to vary. In this case, monthly volumes are then generated with randomisation across different scenarios, all the while ensuring that the conditions for a fixed annual volume and maximum hourly volumes are respected. The conditions of the HPA thus define an envelope in which demand uncertainty is permitted.
A clear distinction is made between in-sample scenarios (scenarios with which the decisions of a particular policy are made), and out-of-sample scenarios (scenarios on which the decisions of a particular policy are tested) [35]. By only evaluating design decisions on their ability to handle new realisations of uncertainty that they have not seen before, we can have confidence that our solutions are not “fluke” design decisions which have over-fit the in-sample set.
@@ 2.2 Objective Function
The risk aversion factor is used to toggle the weighting between the expected value of the operational cost , and the expected value of the worst-case scenario, which is obtained by using the Conditional Value at Risk (CVaR) risk measure [36] with (additional CVAR formulation details provided in the Supplementary Material).
@@ 3.1 Overview
column3 = c,
hline1,7 = -0.08em,
hline2 = -,
Policy & Type Risk
Appetite () No-Resale (NR)
Formulation
1. Deterministic Neutral N/A Yes
2. Deterministic/Rule-based Averse N/A No
3. Stochastic Neutral 0 Yes
4. Stochastic Averse 0.9 No
5. Stochastic Averse 0.9 Yes
@@ 1. Expected Value Problem Policy:
The Expected Value Problem policy is the reference case deterministic risk-neutral policy. It provides a deterministic solution of the model with one input scenario representing an average case for all uncertainties. The no-resale formulation is applied in order to avoid speculative results (Equation (25) is active, as defined in Section 4.3.4). This planning policy is inspired by the prevalence of deterministic modelling in feasibility studies in academia and in industry, as discussed in Section 1.2.1.
@@ 2. Pessimistic Expert Policy:
This policy resembles the previous Expected Value Problem policy, although with the addition of a simple risk-averse heuristic designed to hedge against high spot market prices: the total yearly hydrogen demand (electrical equivalent), is procured using the least-cost PPA options. It is assumed the expert believes diversity to be important, and so acquires 50% of the required yearly production from a low-cost solar PPA and the remaining 50% from a low-cost wind PPA. This calculation is done using the average capacity factor of each park. The remaining design decisions of equipment sizing are then made in a deterministic setting with a single expected value scenario, as for the Expected Value Problem (for uncertain demand this means taking the hour-by-hour average of all in-sample demand scenarios). This policy is inspired by the use of pre-defined hedging ratios in industry as discussed in Section 1.2.3.
@@ 3. Stochastic Risk-Neutral Policy:
This policy uses a risk-neutral 2-stage stochastic formulation. Operational costs of each in-sample scenario have equal weighting (risk aversion weighting in Equation (1) is set to 0). The no-resale formulation is applied (Equation (25) is active). This policy is inspired by risk-neutral stochastic modelling which is typical in capacity expansion models in energy modelling literature as discussed in Section 1.2.2.
@@ 4. Stochastic Risk-Averse Policy:
This policy performs a risk-averse 2-stage stochastic optimization, where the expected value of the operational costs and the operational cost of the worst-case scenario are given weightings of 10% and 90% respectively (risk aversion weighting in Equation (1) is set to 0.9). This policy is inspired by risk-averse modelling in literature with respect to capacity expansion and energy hedging literature as described in Sections 1.2.2 and 1.2.3
@@ 5. Stochastic Risk-Averse No-Resale Policy:
This policy performs a stochastic optimisation with the same risk-aversion weighting as for the preceding policy, but with the no-resale formulation applied (Equation (25) is active). This policy is included to identify the impact of no-resale formulations as used in energy modelling literature as described in Section 1.2.4.
@@ 3.5 Policy Comparison Metrics
The classic metric, theValue of the Stochastic Solution (VSS, [37]) is defined as the difference in performance between a deterministic model, and its stochastic equivalent (typically on expected value). In this study, this metric is applied to both expected value and worst-case LCOH performance, and three other similar metrics are proposed:
@@
The Cost of Demand Uncertainty (CDU), is defined as the proportional difference in performance between solutions optimised for a strict HPA contract (i.e., without demand uncertainty), and those optimised for a flexible contract (i.e., with demand uncertainty).
@@ 4.1 Case Study Selection and Applicability
In this section, a case study of a hydrogen project to supply hydrogen under a HPA in metropolitan France is developed. Specific profiles are used for demand and PPA production, however these may be replaced with any desired profile, making the methodology adjustable to any location with an electricity spot market and hedging instruments comparable to the European market.
In this case study all design decisions are assumed to be continuous within their allowed range. Design decisions could be modelled as mixed integer decisions, which is useful for taking into account incremental equipment sizes or fixed PPA contract capacities. However the objective of this study is to identify the inherent dynamics in the system without imposing exogenous limitations. This approach provides useful information at the initial feasibility and conception stage of project development. When progressing to the detailed planning stage, it would then be appropriate to re-optimise the same model proposed in this paper using mixed integer decisions once equipment suppliers and actual available PPA sizes and costs have been identified.
@@ 4.2.1 Hydrogen Demand
For the case of the Fixed Demand HPA contract, the hydrogen customer does not have the right to vary their future consumption profile. They must obey the original weekday/weekend cyclic profile with no variations week-to-week or month-to-month.
For the Uncertain Demand HPA contracts, the consumer will have some flexibility to vary their future consumption profile (i.e., in response to changing demand for their product). It is assumed that the contract dictates the same annual volume and maximum hourly demand as for the Fixed Demand contract, however the customer will have the right to vary their daily, weekly, and monthly volumes. In scenario generation, these volumes are allowed to vary by +/- 50% with respect to the standard demand scenario used for the strict contract. Three example profiles are shown in Figure 3.
The producer is always obligated to supply. For non-delivery of the required demand, a penalty of €10,000 /MWh is used in in-sample optimisation to strongly discourage planned use of this option, whilst for testing, curtailment is penalised at €1,000 /MWh, which is commonly used as the electrical curtailment cost in other studies [38].
@@ 4.2.2 Hydrogen Equipment
Hydrogen is produced by an electrolyser, followed directly by a compression stage. It may be fed to the customer as it is produced, or be stored and delivered to the customer at a later time. This model uses a simple linear efficiency to model the combined electrolyser + compressor component. This is an approximation of actual electrolysis dynamics which are non-linear, and include discontinuities due to minimum operating levels and distinct operating states [39, 40]. The impact of including a minimum operating point of 25% was evaluated using the Expected Value Problem policy and found to result in no significant change in design decisions, and an increase in final cost of less than 0.1%. It was therefore considered that while modelling these operating states would be preferable, omitting them is not likely to strongly impact the results in this study.
@@ 4.2.3 Day-Ahead Market
Electricity price data for the French day-ahead (spot) market from 2013 to 2022 were obtained [41, 42]. These scenarios represent a range of average annual prices from 30 €/MWh to 275 €/MWh, with a mean of 70 €/MWh. The historical distributions of other metrics were also obtained, including average daily spread, and seasonal price drift.
@@ 4.2.4 Electricity Futures
The first set of instruments that can be used to hedge against high day-ahead prices is energy futures. The products modelled in this study are those with the longest maturity dates in the European market: calendar year (CAL) and quarterly products [41]. The European Energy Exchange (EEX) also offers two categories of delivery profiles: Baseload Futures, for which the quantity of energy bought is delivered in a constant band for every hour across the delivery period; and Peakload Futures, which provides constant power only during the hours 8am to 8pm.
In the context of long-term decision making, even these long-maturity futures are only available up to three years in advance, which is clearly not sufficient for informing investment decisions for project lifetimes of up to 25 years. As such, in the context of this study, design decisions that select these products should be interpreted as a strategy to systematically and progressively hedge consumption during the selected period as soon as these products become available for every year of the project life. They are therefore static global hedging decisions [33].
Risk-neutral futures prices were calculated as the average of the day-ahead prices for the given delivery period in the relevant scenarios. Peakload futures prices are similarly calculated as the expectation of prices within peak hours within the given period. In practice power futures prices involve an uncertain and time-varying risk premium with respect to day-ahead prices [43, 44, 45], however modelling this premium requires making assumptions about market risk factors and their correlation to other variables that are likely to be unreliable and of limited value [33, 34]. Risk-neutral pricing is thus considered to be the most reliable pricing technique for electricity futures for the purposes of this article.
@@ 4.2.5 Power Purchase Agreements (PPA’s)
The second group of energy hedging products considered is a set of renewable PPA’s. A selection of 9 PPA options was included in the model, representing 4 solar photovoltaic (PV) PPA’s and 5 wind PPA’s in different locations in metropolitan France. Data were obtained for 10 years 2013 to 2022 using Renewables Ninja [46, 47].
One method for PPA pricing would consider the future value of energy resale on the spot market, bounded below by the Levelised Cost of Energy (LCOE) of the associated generator [48]. If these two factors were sufficient for predicting PPA prices in the market, PPA prices could be modelled in a risk-neutral manner similarly to electricity futures (possibly including additional risk premiums and option-inspired structures such as in [34] and [49]).
In practice, PPA’s are non-standard and illiquid, and so PPA prices are highly sensitive to difficult-to-model additional factors including but not limited to counterparty risk, negotiating strength and market power, and green regulations [50]. PPA prices that are above and below their expected fair market value are possible (i.e., they cannot be assumed to be risk-neutral with respect to the spot market).
In order to fix PPA prices for the model, first a worst-case (highest cost) LCOE for each site was calculated from the lowest yearly capacity factor (CF), as shown in Table 2. The LCOE’s were used as a guide to establish a rough ordering of PPA prices, which were then chosen arbitrarily to provide a range of reasonably competitive PPA options at levels coherent with recent average PPA prices reported for the European market [51].
@@ 4.2.6 Green Hydrogen Subsidies
Green hydrogen subsidies when applied in this paper are modelled using a requirement for hourly-time matching of hydrogen production with renewable sources (see Section 4.3.3 and [52]). This rule is to be applied in the US and Europe post-2030.
A subsidy level of 3 €/kg is chosen arbitrarily for analysis but is considered reasonable, given the 3 $/kg maximum tax incentive available in the U.S. and the 4.5 €/kg subsidy cap in the European Hydrogen Bank auctions [12, 53] (although it should be noted that the first round of the European auctions cleared much lower, at 0.48 €/kg [54]).
@@ 4.2.7 Cost of Capital
A fixed cost of capital of 5% is used throughout. This is chosen as a reasonable value based on the contextual assumptions of this study — that the project takes place in Europe and with a long-term HPA for 100% of intended production [55]. If merchant risk were involved for hydrogen sales, a larger cost of capital would be appropriate.
@@ 4.2.8 Scenario Sets
Number of hours | ||||
Technical Parameters | ||||
Indices | Electrolyser + Compressor energy | |||
Scenario index | efficiency (MWh H2 / MWh elec) | |||
Hour of the year index | Hydrogen storage input power | |||
Network access time-slot index | efficiency | |||
Futures delivery period index | Hydrogen storage discharge power | |||
@@
Design Decisions | efficiency | |||
Electrolyser power (MW elec) | ||||
Hydrogen storage energy (MWh H2) | Cost Parameters | |||
Hydrogen storage power (MW H2) | Electrolyser CAPEX | |||
Network connection power (MW) | (€/MW elec/year) | |||
Baseload futures, purchased energy | Hydrogen Storage Energy | |||
for period (MWh) | CAPEX (€/MWh H2/year) | |||
Peakload futures, purchased energy | Hydrogen Storage Power | |||
for period (MWh) | CAPEX (€/MW H2/year) | |||
Network access power subscription | Network Connection Power | |||
for time-slot (MW) | CAPEX (€/MW/year) | |||
PPA peak power for park (MWp) | Baseload futures price, | |||
@@
Operational Decisions | Peakload futures price, | |||
Spot market energy bought (MWh) | period (€/MWh) | |||
Spot market energy sold (MWh) | Network power subscription cost | |||
Electrolyser dispatching (MW elec) | for time-slot (€/MW/year) | |||
Hydrogen storage input (MW H2) | PPA energy cost for park (€/MWh) | |||
Hydrogen storage output (MW H2) | Hydrogen demand curtailment | |||
PPA dispatched power, park (MW) | penalty (€/MWh H2) | |||
PPA curtailed power, park (MW) | Green hydrogen subsidy (€/MWh H2) | |||
Hydrogen demand fulfilled (MW H2) | ||||
Hydrogen demand unmet (curtailed) | Design Upper Bounds (UB) | |||
(MW H2) | Electrolyser power UB (MW elec) | |||
Power transferred to the site | Hydrogen storage energy | |||
from the network (MW) | UB (MWh H2) | |||
Subsidy-eligible hydrogen | Hydrogen storage power UB (MW H2) | |||
produced (MWh H2) | Network connection power UB (MW) | |||
@@
Operational States | for period UB (MWh) | |||
Hydrogen storage state-of-charge | Peakload futures, purchased energy | |||
(MWh) | for period UB (MWh) | |||
@@
Operational Uncertainties | for time-slot UB (MW) | |||
Day-ahead market price (€/MWh) | PPA peak power contracted | |||
PPA availability factor | for park UB (MWp) | |||
(% of peak power) | ||||
Hydrogen offtaker demand (MW H2) | Objective Function Paramaters | |||
@@
Risk aversion weighting {[}0:1{]} |

### v3

@@ 1.4 Contributions
An energy market focused 2-stage stochastic model of an electrolytic hydrogen production plant in the context of an HPA, with co-optimisation of plant capacities, multiple electricity sourcing options, and the possibility of green hydrogen subsidies.
@@
Hedging energy resale (resale-enabled or no-resale formulations).
@@ 2.1 Overview
HPA’s in practice may have different conditions that are designed to reduce demand uncertainty, including but not limited to minimum and maximum yearly, monthly, daily, and hourly volumes. In this study, an HPA with fixed demand corresponds to a contract where the consumer is bound to follow precisely their original profile (i.e., all volumes down to hourly volumes are fixed, and no randomisation is included). Uncertainty can then be added by relaxing some of these rules, for example, by allowing monthly volumes to vary. In this case, monthly volumes are then generated with randomisation across different scenarios, all the while ensuring that the conditions for a fixed annual volume and maximum hourly volumes are respected. The conditions of the HPA thus define an envelope in which demand uncertainty is permitted.
A clear distinction is made between in-sample scenarios (scenarios with which the decisions of a particular policy are made), and out-of-sample scenarios (scenarios on which the decisions of a particular policy are tested) (King2012ModelingProgramming). By only evaluating design decisions on their ability to handle new realisations of uncertainty that they have not seen before, we can have confidence that our solutions are not “fluke” design decisions which have over-fit the in-sample set.
@@ 2.2 Objective Function
The risk aversion factor is used to toggle the weighting between the expected value of the operational cost , and the expected value of the worst-case scenario, which is obtained by using the Conditional Value at Risk (CVaR) risk measure (Rockafellar2000OptimizationValue-at-risk) with (additional CVAR formulation details provided in the Supplementary Material).
@@ 3.1 Overview
| Policy | Type | Risk | ||
| Appetite | () | No-Resale (NR) | ||
| Formulation | ||||
| 1. | Deterministic | Neutral | N/A | Yes |
| 2. | Deterministic/Rule-based | Averse | N/A | No |
| 3. | Stochastic | Neutral | 0 | Yes |
| 4. | Stochastic | Averse | 0.9 | No |
| 5. | Stochastic | Averse | 0.9 | Yes |
@@ 1. Expected Value Problem Policy:
The Expected Value Problem policy is the reference case deterministic risk-neutral policy, with co-optimised hedging and equipment capacity decisions. It provides a deterministic solution of the model with one input scenario representing an average case for all uncertainties. The no-resale formulation is applied in order to avoid speculative results (Equation (25) is active, as defined in Section 4.3.4). This planning policy is inspired by the prevalence of deterministic modelling in feasibility studies in academia and in industry, as discussed in Section 1.2.1.
@@ 2. Pessimistic Expert Policy:
This policy resembles the previous Expected Value Problem policy, although without equipment and hedging co-optimisation. For the hedging decisions, a simple exogenous risk-averse heuristic is used to hedge against high spot market prices: the total yearly hydrogen demand (electrical equivalent), is procured using the least-cost PPA options. It is assumed the expert believes diversity to be important, and so acquires 50% of the required yearly production from a low-cost solar PPA and the remaining 50% from a low-cost wind PPA. This calculation is done using the average capacity factor of each park. The remaining design decisions of equipment sizing are then made in a deterministic setting with a single expected value scenario, as for the Expected Value Problem (for uncertain demand this means taking the hour-by-hour average of all in-sample demand scenarios). This policy is inspired by separately optimised hedging decisions and the use of pre-defined hedging ratios, which are in industry as discussed in Section 1.2.3.
@@ 3. Stochastic Risk-Neutral Policy:
This policy uses a risk-neutral 2-stage stochastic formulation, with co-optimised hedging and equipment capacity decisions. Operational costs of each in-sample scenario have equal weighting (risk aversion weighting in Equation (1) is set to 0). The no-resale formulation is applied (Equation (25) is active). This policy is inspired by risk-neutral stochastic modelling which is typical in capacity expansion models in energy modelling literature as discussed in Section 1.2.2.
@@ 4. Stochastic Risk-Averse Policy:
This policy performs a co-optimised risk-averse 2-stage stochastic optimisation, where the expected value of the operational costs and the operational cost of the worst-case scenario are given weightings of 10% and 90% respectively (risk aversion weighting in Equation (1) is set to 0.9). This policy is inspired by risk-averse modelling in literature with respect to capacity expansion and energy hedging literature as described in Sections 1.2.2 and 1.2.3.
While adjustment of either or the CVAR can achieve similar risk-averse solutions, the main consideration is the use of settings that provide a high level of risk aversion without being fully robust (as this ignores many scenarios in the in-sample set and tends to overfit the worst-case scenario).
@@ 5. Stochastic Risk-Averse No-Resale Policy:
This policy performs again a co-optimised stochastic optimisation with the same risk-aversion weighting as for the preceding policy, but with the no-resale formulation applied (Equation (25) is active). This policy is included to identify the impact of no-resale formulations as used in energy modelling literature as described in Section 1.3.2.
@@ 3.5 Policy Comparison Metrics
The classic metric, the Value of the Stochastic Solution, or VSS, (Birge2010IntroductionProgramming) is defined as the difference in performance between a deterministic model, and its stochastic equivalent (typically on expected value). In this study, this metric is applied to both expected value and worst-case LCOH performance, and three other similar metrics are proposed:
@@
The Cost of Demand Uncertainty (CDU), is defined as the proportional difference in performance between solutions obtained using the same planning policy, but for a Fixed Demand HPA contract, and an Uncertain Demand HPA contract.
@@ 4.1 Case Study Selection and Applicability
In this section, a case study of a hydrogen project to supply hydrogen under an HPA in metropolitan France is developed. Specific profiles are used for demand and PPA production, however these may be replaced with any desired profile, making the methodology adjustable to any location with an electricity spot market and hedging instruments comparable to the European market.
In this case study all design decisions are assumed to be continuous within their allowed range, including both hedging decisions and equipment capacity decisions. Design decisions could be modelled as mixed integer decisions, which is useful for taking into account incremental equipment sizes or fixed PPA contract capacities. However the objective of this study is to identify the inherent dynamics in the system without imposing exogenous limitations. This approach provides useful information at the initial feasibility and conception stage of project development. When progressing to the detailed planning stage, it would then be appropriate to re-optimise the same model proposed in this paper using mixed integer decisions once equipment suppliers and actual available PPA sizes and costs have been identified.
Previous studies have indicated that the inclusion of lithium-ion batteries is not currently cost-effective for the sole purpose of time-shifting electrical demand of the electrolyser (Superchi2023Techno-economicSteelmaking; Palmer2023Risk-conscious.), however the inclusion of value stacking with other services may change this result. Considering additional revenue sources is beyond the scope of this study.
@@ 4.2.1 Hydrogen Demand
For the case of the Fixed Demand HPA contract, the hydrogen customer does not have the right to vary their future consumption profile. They must obey the original weekday/weekend cyclic profile (see Figure 3) with no variations week-to-week or month-to-month.
For the Uncertain Demand HPA contracts, the consumer will have some flexibility to vary their future consumption profile (i.e., in response to changing demand for their product). It is assumed that the contract dictates the same annual volume and maximum hourly demand as for the Fixed Demand contract, however the customer will have the right to vary their daily, weekly, and monthly volumes. In scenario generation, these volumes are allowed to vary by +/- 50% with respect to the standard demand scenario used for the strict contract. Three example scenarios are shown in Figure 4, and a comparative illustration of their associated month-to-month volume changes is provided in the Supplementary Material.
Another important condition to be defined in the HPA is a monetary penalty for unfulfilled (curtailed) demand. In our case study, for non-delivery of the required demand for a given hour, a penalty of €10,000 /MWh is used in in-sample optimisation to strongly discourage planned use of this option, whilst for testing, curtailment is penalised at €1,000 /MWh, which is commonly used as the electrical curtailment cost in other studies (Jodry2023IndustrialOptimisation).
@@ 4.2.2 Hydrogen Equipment
Hydrogen is produced by an electrolyser, followed directly by a compression stage. It may be fed to the customer as it is produced, or be stored and delivered to the customer at a later time. This model uses a simple linear efficiency to model the combined electrolyser + compressor component. This is an approximation of actual electrolysis dynamics which are non-linear, and include discontinuities due to minimum operating levels and distinct operating states (Zheng2022OptimalDynamics; Baumhof2023OptimizationNecessary). The impact of including an upper-range minimum operating point of 25% (Baumhof2023OptimizationNecessary) was evaluated using the Expected Value Problem policy and found to result in no significant change in design decisions, and an increase in final cost of less than 0.1%. It was therefore considered that while modelling these operating states would be preferable, omitting them is not likely to strongly impact the results in this study.
@@ 4.2.3 Day-Ahead Market
Electricity price data for the French day-ahead (spot) market from 2013 to 2022 were obtained (EEX2022EuropeanEEX; ENTSOe2022ENTSOePlatform). These scenarios represent a range of average annual prices from 30 €/MWh to 275 €/MWh, with a mean of 70 €/MWh. The historical distributions of other metrics were also obtained, including average daily spread, and seasonal price drift.
@@ 4.2.4 Electricity Futures
The first set of instruments that can be used to hedge against high day-ahead prices is energy futures. The products modelled in this study are those with the longest maturity dates in the European market: calendar year (CAL) and quarterly products (EEX2022EuropeanEEX). The European Energy Exchange (EEX) also offers two categories of delivery profiles: Baseload Futures, for which the quantity of energy bought is delivered in a constant band for every hour across the delivery period; and Peakload Futures, which provides constant power only during the hours 8am to 8pm.
In the context of long-term decision making, even these long-maturity futures are only available up to three years in advance, which is clearly not sufficient for informing investment decisions for project lifetimes of up to 25 years. As such, in the context of this study, design decisions that select these products should be interpreted as a strategy to systematically and progressively hedge consumption during the selected period as soon as these products become available for every year of the project life. They are therefore static global hedging decisions (Dimoski2023DynamicRisks).
Risk-neutral futures prices were calculated as the average of the day-ahead prices for the given delivery period in the relevant scenarios. Peakload futures prices are similarly calculated as the expectation of prices within peak hours within the given period. In practice power futures prices involve an uncertain and time-varying risk premium with respect to day-ahead prices (Benth2013AnMarkets; Oliveira2021AnalysisAversion; Secomandi2022QuadraticOperations), however modelling this premium requires making assumptions about market risk factors and their correlation to other variables that are likely to be unreliable and of limited value (Tranberg2020ManagingAgreements; Dimoski2023DynamicRisks). Risk-neutral pricing is thus considered to be the most reliable pricing technique for electricity futures for the purposes of this article.
@@ 4.2.5 Power Purchase Agreements (PPA’s)
The second group of energy hedging products considered is a set of renewable PPA’s. A selection of 9 PPA options was included in the model, representing 4 solar photovoltaic (PV) PPA’s and 5 wind PPA’s in different locations in metropolitan France. Data were obtained for 10 years 2013 to 2022 using Renewables Ninja (Pfenninger2016Long-termData; Staffell2016UsingOutput).
One method for PPA pricing would consider the future value of energy resale on the spot market, bounded below by the Levelised Cost of Energy (LCOE) of the associated generator (Qorbanian2024ValuationProcurement). If these two factors were sufficient for predicting PPA prices in the market, PPA prices could be modelled in a risk-neutral manner similarly to electricity futures (possibly including additional risk premiums and option-inspired structures such as in Tranberg2020ManagingAgreements and Pombo-Romero2024AssessingPPAs).
In practice, PPA’s are non-standard and illiquid, and so PPA prices are highly sensitive to difficult-to-model additional factors including but not limited to counterparty risk, negotiating strength and market power, and green regulations (Mili2025GreenAgreements). PPA prices that are above and below their expected fair market value are possible (i.e., they cannot be assumed to be risk-neutral with respect to the spot market).
In order to fix PPA prices for the model, first a worst-case (highest cost) LCOE for each site was calculated from the lowest yearly capacity factor (CF), as shown in Table 2. The LCOE’s were used as a guide to establish a rough ordering of PPA prices, which were then chosen arbitrarily to provide a range of reasonably competitive PPA options at levels coherent with recent average PPA prices reported for the European market (Joinville2023European3Q).
@@ 4.2.6 Green Hydrogen Subsidies
Green hydrogen subsidies when applied in this paper are modelled using a requirement for hourly-time matching of hydrogen production with renewable sources. This rule is to be applied in the US and Europe post-2030 (Guillotin2025HydrogenComparison).
A subsidy level of 3 €/kg is chosen arbitrarily for analysis but is considered reasonable, given the 3 $/kg maximum tax incentive available in the U.S. and the 4.5 €/kg subsidy cap in the European Hydrogen Bank auctions (EuropeanCommission2023CommissionProduction.; USDepartmentofEnergy2023AssessingCredit), although it should be noted that the first round of the European auctions cleared much lower, at 0.48 €/kg (EuropeanCommission2024EuropeanEurope).
@@ 4.2.7 Cost of Capital
A fixed cost of capital of 5% is used throughout. This is chosen as a reasonable value based on the contextual assumptions of this study — that the project takes place in Europe and with a long-term HPA for 100% of intended production (InternationalEnergyAgency2021TheTransitions). If merchant risk were involved for hydrogen sales, a larger cost of capital would be appropriate.
@@ 4.2.8 Scenario Sets
| Number of hours | Technical Parameters | |||
| Electrolyser + Compressor energy | ||||
| Indices | efficiency (MWh H2 / MWh elec) | |||
| Scenario index | ||||
| Hydrogen storage input power | ||||
| Hour of the year index | efficiency | |||
| Network access time-slot index | Hydrogen storage discharge power | |||
| efficiency | ||||
| Futures delivery period index | ||||
| Hydrogen storage energy | ||||
@@
| Design Decisions | Cost Parameters | |||
| Electrolyser power (MW elec) | Electrolyser CAPEX | |||
| (€/MW elec/year) | ||||
| Hydrogen storage energy (MWh H2) | ||||
@@
| Hydrogen storage power (MW H2) | CAPEX (€/MWh H2/year) | |||
| Network connection power (MW) | Hydrogen Storage Power | |||
| CAPEX (€/MW H2/year) | ||||
| Baseload futures, purchased energy | ||||
| for period (MWh) | Network Connection Power | |||
| CAPEX (€/MW/year) | ||||
| Peakload futures, purchased energy | ||||
| for period (MWh) | Baseload futures price, | |||
@@
| Network access power subscription | ||||
| for time-slot (MW) | Peakload futures price, | |||
| period (€/MWh) | ||||
| PPA peak power for park (MWp) | ||||
| Network power subscription cost | ||||
| Operational Decisions | for time-slot (€/MW/year) | |||
| Spot market energy bought (MWh) | ||||
| PPA energy cost for park (€/MWh) | ||||
| Spot market energy sold (MWh) | ||||
| Hydrogen demand curtailment | ||||
| Electrolyser dispatching (MW elec) | penalty (€/MWh H2) | |||
| Hydrogen storage input (MW H2) | Green hydrogen subsidy (€/MWh H2) | |||
| Hydrogen storage output (MW H2) | Design Upper Bounds (UB) | |||
| Electrolyser power UB (MW elec) | ||||
| PPA dispatched power, park (MW) | ||||
| Hydrogen storage energy | ||||
| PPA curtailed power, park (MW) | UB (MWh H2) | |||
| Hydrogen demand fulfilled (MW H2) | Hydrogen storage power UB (MW H2) | |||
| Hydrogen demand unmet (curtailed) | Network connection power UB (MW) | |||
| (MW H2) | ||||
@@
| Power transferred to the site | for period UB (MWh) | |||
| from the network (MW) | ||||
| Peakload futures, purchased energy | ||||
| Subsidy-eligible hydrogen | for period UB (MWh) | |||
| produced (MWh H2) | ||||
@@
| Operational States | for time-slot UB (MW) | |||
| Hydrogen storage state-of-charge | ||||
| (MWh) | PPA peak power contracted | |||
| for park UB (MWp) | ||||
| Operational Uncertainties | ||||
| Day-ahead market price (€/MWh) | Objective Function Paramaters | |||
@@
| PPA availability factor | ||||
| (% of peak power) | Risk aversion weighting {{[}}0:1{{]}} | |||
| Hydrogen offtaker demand (MW H2) | ||||